OpenStudy (anonymous):

convert to a fraction 0,081081081

OpenStudy (anonymous):

\[\frac{ repeating digits }{ 10^{number of repeating digits} }\]

OpenStudy (anonymous):

then simplify your fraction..

OpenStudy (anonymous):

do it @jesi1201

OpenStudy (anonymous):

sorry, i forgot that its a repeating decimal... i first thought of a terminating decimal.. my bad..

OpenStudy (anonymous):

can u do it .. if so then plz do it ..

OpenStudy (anonymous):

is the given 0.081081081... right? with 081 repeating?

OpenStudy (anonymous):

ok then

OpenStudy (anonymous):

081 block is repeating after decimal point .. how to solve

OpenStudy (anonymous):

if it's a repeating decimal, let n=0.081081... (eq. 1) multiply both sides by 10^3, since there are 3 repeating digits... we have, 1000n = 081.081081... (eq. 2) subtract eq 1 from eq 2, that is 1000n = 081.081081... - n = 0.081081... ___________________________________ 999n = 81 \[n = \frac{ 81 }{ 999}\] simplify...

OpenStudy (anonymous):

well i have another idea .. to multiply it by 1000 and subtracting the same number from result will make it definite number i think wat do u say @jesi1201

OpenStudy (anonymous):

oh ya u did the same

OpenStudy (anonymous):

ya its done thank u @jesi1201 u made some thing fresh to me that i learned 8 years ago :P

OpenStudy (anonymous):

thanks, sorry again for my mistake earlier...

OpenStudy (anonymous):

no its ok . actuallly i did nt get that mistake so i asked u